OpenStudy (anonymous):

I'm think it's b is it? Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form. <1, 1>

geerky42 (geerky42):

\(Unit~Vector = \dfrac{\mathbf u}{||~\mathbf u~||}\)

OpenStudy (anonymous):

So how do I find the denominator?

geerky42 (geerky42):

\(||~\mathbf u~||\) is magnitude of \(\mathbf u\), So you basically use pythagorean theorem to find it.  \[||~\mathbf u~|| = \sqrt{(-4)^2+(-3)^2}\] Does that make sense?

OpenStudy (anonymous):

Umm let me try to solve it

OpenStudy (anonymous):

So 5i

OpenStudy (anonymous):

Or would the denominator be 25 and not five @geerky42

geerky42 (geerky42):

\(=\sqrt{(-4)^2+(-3)^2}\\~\\=\sqrt{16+9} \\~\\=\sqrt{25}\\~\\=5\) So denominator would be 5.

geerky42 (geerky42):

So you have unit vector = \(\dfrac{\mathbf u}{5}\)

OpenStudy (anonymous):

Oh so if you sqaure a number it becomes positive

geerky42 (geerky42):

Yes, squaring negative number will get you positive number.

geerky42 (geerky42):

Always.

OpenStudy (anonymous):

Oh okay thanks :)

geerky42 (geerky42):

Welcome! You can handle the rest on your own?

OpenStudy (anonymous):

Umm I believe so but I have another question that's a tad different Two forces with magnitudes of 100 and 50 pounds act on an object at angles of 50° and 160° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

Two forces with magnitudes of 100 and 50 pounds act on an object at angles of 50° and 160° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

Pounds act on an object at angles of 50° and 160° respectively.

OpenStudy (anonymous):

Degree in place of all those boxes

OpenStudy (anonymous):

@geerky42